Question 1, 2 And 3

A = [ 2 5 ] is the given matrix.
       [ 3 5 ]

1) (2marks) encode the message "TEST".

2)(2marks) find the inverse of A. Note that your matrix entries should be integers in the range 0,...25.

3) (2marks) use your answer above to decode the message "ETAJ"
Answer #1

Sol 2 5 Given A = L3 5 Given message TEST . Numbro alphabets A, B .... 2 with 0,1,.... 25 T E s T 19 4 18 19 Take two numbers2 Al Adja det A 2 5 A = AdjA= 5.-5 A 3. 5 -3 2 deta=10-15=-5A multiplicative inverse of a modo is the integer y such that XY=lmos we Get detA==5 ir - =-5% 5 mod 26 26= so deta 5 -(5-modSo A L AdjA= 31 Iss is 5 -5 -3 2 deta -93 Oradea AV= /25 ī . ll 3 Given 4 9 ole coding the ЕТА A J 19.0 the Message E А -> 12

Plz do Upvote